OpenStudy (anonymous):

If c is the number that satisfies the conclusion of the Mean Value Theorem for f(x)=x^3-2x^2 on the interval 0<=x<=2, find c.

OpenStudy (anonymous):

\[ f'(c) = \frac {f(2)-f(0)}{2-0} \]

OpenStudy (anonymous):

But I got 0 for that.

OpenStudy (anonymous):

I don't get it. The answer should be 4/3.

OpenStudy (anonymous):

Sorry, hang on

OpenStudy (anonymous):

\[ f'(x)=3 x^2-4 x=0\\ x(3 x -4)=0\\ x=0 \\ x=\frac 4 3 \]

OpenStudy (anonymous):

Get it. Thanks.

OpenStudy (anonymous):

YW